OpenStudy (anonymous):

h y varies directly as the square of x. When x = 1, y = 3. What is the value of y when x = 3?

OpenStudy (anonymous):

y = kx^2 3 = k(1)^2, k = 3 so when x = 3, using y = 3x^2 y = 27